# Bayesian Inference Under Differential Privacy: Prior Selection Considerations with Application to Univariate Gaussian Data and Regression

### v1

@@ Appendix E Constraints for the Univariate Gaussian Setting
Let $Y_{1},\ldots,Y_{n}\sim{\mathcal{N}}(\mu,\sigma^{2})$ and suppose each $Y_{i}\in[0,1]$. If $\mu$ is known, then $\sigma^{2}\in[0,\mu(1-\mu)]$. If $\sigma^{2}$ is known, then $\mu\in[1/2-\sqrt{1/4-\sigma^{2}},1/2+\sqrt{1/4-\sigma^{2}}]$.
@@
If $Y_{1},\ldots,Y_{n}\sim{\mathcal{N}}(\mu,\sigma^{2})$ and $Y_{i}\in[0,1]$, then $\sigma^{2}\leq 1/4$.
@@
Let $Y_{1},\ldots,Y_{n}\sim{\mathcal{N}}(\mu,\sigma^{2})$ and suppose each $Y_{i}\in[0,1]$. If $\bar{Y}$ is known, then $S^{2}\in[0,n/(n-1)\bar{Y}(1-\bar{Y})]$. If $S^{2}$ is known, then $\bar{Y}\in[1/2-\sqrt{1/4-(n-1)/n\cdot S^{2}},1/2+\sqrt{1/4-(n-1)/n\cdot S^{2}}]$.
@@ Appendix F Determining Whether the Posterior Distribution is Proper
For confidential data $Y_{1},\ldots,Y_{n}\overset{iid}{\sim}{\mathcal{N}}(\mu,\sigma^{2})$ where sufficient statistics $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$ and $S^{2*}\sim\textsf{Lap}(S^{2},1/(\varepsilon_{2}n))$ are released, if an analyst has prior $p(\mu,\sigma^{2})\propto(\sigma^{2})^{-1}$, then their posterior $p(\mu,\sigma^{2}\mid\bar{Y}^{*},S^{2*})$ is not a proper probability distribution.
@@
For confidential data $Y_{1},\ldots,Y_{n}\overset{iid}{\sim}{\mathcal{N}}(\mu,\sigma^{2})$ where $n>3$ and sufficient statistics $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$ and $S^{2*}\sim\textsf{Lap}(S^{2},1/(\varepsilon_{2}n))$ are released, if an analyst has prior $p(\mu,\sigma^{2})\propto 1$, then their posterior $p(\mu,\sigma^{2}\mid\bar{Y}^{*},S^{2*})$ is a proper probability distribution.
@@
By (78, the double integral in the second bracket is equal to 1. Thus, the quantity of interest reduces to
@@ Appendix H Additional Simulation Studies
Figure 8 presents the RMSE from estimating $\mu$ with the posterior mode in the simulation study from Section 4 used to make Figure 2 of the main text. The RMSE is uniformly lower in the constrained analysis than in the unconstrained analysis. For the constrained analysis, a regression of log RMSE on log $n$ has a slope of $-1$ (with $R^{2}>0.999$), indicating that RMSE decays proportionally to $1/n$
@@
Figure 9: Empirical coverage rate of 95% HPD intervals for $\mu$ (top), average 95% HPD interval length for $\mu$ (middle), and root mean square error for estimating $\mu$ (bottom) for different sample sizes. Results based on 10,000 simulated datasets $Y_{i}\in[0,1]$ released with $\varepsilon_{1}=\varepsilon_{2}=1$ and analyzed with prior $p(\mu,\sigma^{2})\propto 1$. Analyses with constraints accounted for are solid lines and not accounted for are dashed lines. Data generating model is either ${\mathcal{N}}(\mu=0.1,\sigma^{2}=0.04^{2})$ (red) or ${\mathcal{N}}(\mu=0.5,\sigma^{2}=0.2^{2})$ (blue). Each Gibbs sampler is run for 20,000 iterations.
Appendix I Constraints in Simple Linear Regression
@@
In this section, we derive constraints on the model proposed by [5] in the simple linear regression setting. Letting ${\mathbf{X}}=[{\mathbf{1}}_{n}\hskip 5.78172pt{\mathbf{x}}_{1}]$, the authors release ${\mathbf{X}}^{\top}{\mathbf{X}}$, ${\mathbf{X}}^{\top}{\mathbf{Y}}$, and ${\mathbf{Y}}^{\top}{\mathbf{Y}}$, which include the following elements.
${\mathbf{X}}^{\top}{\mathbf{X}}=\begin{bmatrix}{\mathbf{1}}_{n}^{\top}{\mathbf{1}}_{n}&{\mathbf{1}}^{\top}{\mathbf{x}}_{1}\\ {\mathbf{x}}_{1}^{\top}{\mathbf{1}}_{n}&{\mathbf{x}}_{1}^{\top}{\mathbf{x}}_{1}\end{bmatrix},\qquad{\mathbf{X}}^{\top}{\mathbf{Y}}=\begin{bmatrix}{\mathbf{1}}_{n}^{\top}{\mathbf{Y}}\\ {\mathbf{x}}_{1}^{\top}{\mathbf{Y}}\end{bmatrix}.$ (106)
@@
    Since $x_{i}\leq 1$ for all $i$, it follows that ${\mathbf{x}}_{1}^{\top}{\mathbf{1}}_{n}={\mathbf{1}}^{\top}{\mathbf{x}}_{1}=\sum_{i=1}^{n}x_{i}\leq n$.
@@
Appendix J Compute Details
In this section, we describe information on the computer resources needed to reproduce the experiments. We note that Figures 1, 3, 4, and 6 can be run on a personal computer in a few seconds and do not require any additional computing resources. Figure 5 can be run on a personal computer in under five minutes and so also does not require any additional computing resources.

### v2

@@ Appendix E CONSTRAINTS FOR THE UNIVARIATE GAUSSIAN SETTING
Let $Y_{i}\in[0,1]$ have moments $E[Y_{i}]=\mu$ and $V[Y_{i}]=\sigma^{2}$. It follows that $\sigma^{2}\in[0,\mu(1-\mu)]$ and $\mu\in[1/2-\sqrt{1/4-\sigma^{2}},1/2+\sqrt{1/4-\sigma^{2}}]$.
@@
If $Y_{i}\in[0,1]$, then $V[Y_{i}]=\sigma^{2}\leq 1/4$.
@@
Let $Y_{1},\ldots,Y_{n}$ be such that each $Y_{i}\in[0,1]$. Then $S^{2}\in[0,n/(n-1)\bar{Y}(1-\bar{Y})]$ and $\bar{Y}\in[1/2-\sqrt{1/4-(n-1)/n\cdot S^{2}},1/2+\sqrt{1/4-(n-1)/n\cdot S^{2}}]$.
@@ Appendix F DETERMINING WHETHER THE POSTERIOR DISTRIBUTION IS PROPER
For confidential data $Y_{1},\ldots,Y_{n}\overset{iid}{\sim}{\mathcal{N}}(\mu,\sigma^{2})$ where $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$ and $S^{2*}\sim\textsf{Lap}(S^{2},1/(\varepsilon_{2}n))$ are released, if an analyst has prior $p(\mu,\sigma^{2})\propto(\sigma^{2})^{-1}$, then their posterior $p(\mu,\sigma^{2}\mid\bar{Y}^{*},S^{2*})$ is not a proper probability distribution.
@@
For confidential data $Y_{1},\ldots,Y_{n}\overset{iid}{\sim}{\mathcal{N}}(\mu,\sigma^{2})$ where $n>3$ and $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$ and $S^{2*}\sim\textsf{Lap}(S^{2},1/(\varepsilon_{2}n))$ are released, if an analyst has prior $p(\mu,\sigma^{2})\propto 1$, then their posterior $p(\mu,\sigma^{2}\mid\bar{Y}^{*},S^{2*})$ is a proper probability distribution.
@@
By (77), the double integral in the second bracket is equal to 1. Thus, the quantity of interest reduces to
@@ Appendix H ADDITIONAL SIMULATION STUDIES
Figure 8 presents the RMSE from estimating $\mu$ with the posterior mode in the simulation study from Section 4 used to make Figure 2 of the main text. The RMSE is uniformly lower in the constrained analysis than in the unconstrained analysis. For the unconstrained analysis, a regression of log RMSE on log $n$ has a slope of $-1$ (with $R^{2}>0.999$), indicating that RMSE decays proportionally to $1/n$
@@
Figure 9: Average 95% HPD interval length for $\mu$ (top), root mean square error for estimating $\mu$ (middle), and empirical coverage rate of 95% HPD intervals for $\mu$ (bottom) for different sample sizes. Results based on 10,000 simulated datasets $Y_{i}\in[0,1]$ released with $\varepsilon_{1}=\varepsilon_{2}=1$ and analyzed with prior $p(\mu,\sigma^{2})\propto 1$. Analyses with constraints accounted for are solid lines and not accounted for are dashed lines. Data generating model is either ${\mathcal{N}}(\mu=0.1,\sigma^{2}=0.04^{2})$ (red) or ${\mathcal{N}}(\mu=0.5,\sigma^{2}=0.2^{2})$ (blue). Each Gibbs sampler is run for 20,000 iterations.
Appendix I CONSTRAINTS COMPARISON
In the main text, we focus on accounting for bounds on the data values via constraints on the sample moments (and the model parameters). An alternative method is to enforce the constraint on the data values themselves via the likelihood function instead of via the sample moments (while still constraining the parameters to the feasible region determined by Theorem 2). In this section, we compare inferences from this alternative method of accounting for the constraints to inferences from our approach.
The Gibbs samplers in Section 2.2, as well as methods based on sufficient statistics like the samplers of Bernstein & Sheldon (2018, 2019), are unable to easily incorporate constraints on likelihoods. Thus, to make this comparison, we develop and utilize a constraint-aware version of the Gibbs sampler of Ju et al. (2022).
I.1 Enforcing Constraints on the Likelihood
To begin, we describe the Gibbs sampler proposed by Ju et al. (2022) for the univariate Gaussian setting where we do not have constraints on data values. At the $t^{\textrm{th}}$ iteration, we sample from $n+2$ full conditionals: those of $\mu$, $\sigma^{2}$, and each of the confidential $Y_{i}$. The full conditionals for $\mu$ and $\sigma^{2}$ are identical to the public setting (and so have the forms in Appendix D). For each $Y_{i}$, a proposal is drawn from $Y_{i}^{\prime}\sim{\mathcal{N}}(\mu_{(t)},\sigma^{2}_{(t)})$. Let $\bar{Y}_{\textsf{prev}}$ and $\bar{Y}^{\prime}$ be the sample mean with and without the proposal, respectively, i.e.,
$\bar{Y}_{\textsf{prev}}=\frac{1}{n}\left(\sum_{j=1}^{i-1}Y_{i(t)}+Y_{i(t-1)}+\sum_{j=i+1}^{n}Y_{i(t-1)}\right),\qquad\bar{Y}^{\prime}=\frac{1}{n}\left(\sum_{j=1}^{i-1}Y_{i(t)}+Y_{i}^{\prime}+\sum_{j=i+1}^{n}Y_{i(t-1)}\right).$ (105)
Let $S^{2}_{\textsf{prev}}$ and $S^{2\prime}$ be defined similarly for the sample variance. Notably, for a given proposal $Y_{i}^{\prime}$, the updated moments $(\bar{Y}^{\prime},S^{2\prime})$ can be computed from $(\bar{Y}_{\textsf{prev}},S^{2}_{\textsf{prev}})$ in constant time. Then, the proposed $Y_{i}$ is accepted with probability,
$r=\min\left\{1,\exp\left[-\varepsilon_{1}n\left(|\bar{Y}^{*}-\bar{Y}^{\prime}|-|\bar{Y}^{*}-\bar{Y}_{\textsf{prev}}|\right)-\varepsilon_{2}n\left(|S^{2*}-S^{2\prime}|-|S^{2*}-S^{2}_{\textsf{prev}}|\right)\right]\right\}.$ (106)
One can verify empirically that this Gibbs sampler produces a posterior distribution identical to that of our proposed Gibbs sampler without the constraints on the sample moments. Nonetheless, as described in Section 2.2, the sampler of Ju et al. (2022) has computational complexity $\mathcal{O}(nT)$, where $T$ is the number of iterations of the Gibbs sampler, whereas our proposed sampler has computational complexity $\mathcal{O}(T)$.
We now describe how to modify the sampler of Ju et al. (2022) to enforce the constraint that each $Y_{i}\in[a,b]$. Basically, we apply methods similar to those described in Section 3. For $\mu$ and $\sigma^{2}$, we constrain the full conditionals as in Theorem 2. We also specify a proper prior distribution, which is a required assumption to utilize the sampler of Ju et al. (2022). For the illustrations below, we utilize the proper prior $p(\mu,\sigma^{2})\propto 1$ over the feasible region of the parameter space. For each $Y_{i}$, we enforce the constraint on the likelihood as part of the sampler by rejecting any proposed draws outside of $[a,b]$.
I.2 Comparing the Constrained Posterior Distributions
Consider a setting with sample size $n=50$ and privacy budget $\varepsilon_{1}=\varepsilon_{2}=0.25$ where it is known that $Y_{i}\in[0,1]$ for all $i$. As a first illustration, we suppose that $\bar{Y}=0.5$ is in the center of the constrained region and $S^{2}=0.2^{2}$ is fairly large. Figure 10 plots the posterior distribution of $(\mu,\sigma^{2}\mid\bar{Y}^{*},S^{2*})$ for a particular draw of $(\bar{Y}^{*},S^{2*})$ from the Laplace mechanism. As shown in the left panel of Figure 10, our proposed Gibbs sampler from Section 2.2 yields a posterior distribution with mode around the released values. The posterior has considerable uncertainty and places probability density throughout the feasible region. As shown in the right panel of Figure 10, the constrained sampler described above in Section I.1 also yields a posterior with its estimate for $\mu$ close to the released $\bar{Y}^{*}$. However, its estimate for $\sigma^{2}$ is considerably smaller than the released $S^{2*}$. The posterior distribution has substantially less uncertainty, concentrating the probability density in the region where $\sigma^{2}$ is small and $\mu$ is near $\bar{Y}^{*}$.
The difference in the two posterior distributions stems from the implications of constructing draws of the sample moments $(\bar{Y},S^{2})$ by individually sampling constrained $Y_{i}$, as in our adaptation of the sampler of Ju et al. (2022), or sampling directly from the sample moments’ full conditional distributions, as in our proposed approach. Recall from Theorem 3 that any data bounded in $[a,b]$ has bounded sample variance, $S^{2}$. Since the modified sampler of Ju et al. (2022) favors values of $Y_{i}$ that have high density under a constrained normal model, i.e, are far from the bounds $a$ and $b$, the $Y_{i}$ have a sample variance in the lower part of the feasible range for $S^{2}$. As a result, the modified sampler of Ju et al. (2022) favors relatively small values of $\sigma^{2}$, as observed in Figure 10. In contrast, our sampler gives posterior density to values of $S^{2}$ in the upper part of the feasible range indicated by the bound in Theorem 3, as it is not tied to finding Gaussian models that allow for high density samples of each individual $Y_{i}$.
Figure 10: The joint posterior distribution for $(\mu,\sigma^{2})$ under a uniform prior with $n=50$, $\varepsilon_{1}=\varepsilon_{2}=0.25$, $a=0$, and $b=1$. The unreleased confidential statistics $(\bar{Y}=0.5,S^{2}=0.2^{2})$ are represented by the green triangle, the released statistics $(\bar{Y}^{*}=0.43,S^{2}=0.28^{2})$ are represented by the red circle, and the analyst’s posterior mode is represented by the blue diamond. The left and right panels provide the posterior when the statistics are constrained and when the likelihood is constrained, respectively. The shaded area represents the feasible region for $(\mu,\sigma^{2})$ from Theorem 2. This plot is based on 25,000 Gibbs iterations, pruned to every 5th draw.
Figure 11: The joint posterior distribution for $(\mu,\sigma^{2})$ under a uniform prior with $n=50$, $\varepsilon_{1}=\varepsilon_{2}=0.25$, $a=0$, and $b=25$. The unreleased confidential statistics $(\bar{Y}=0.5,S^{2}=0.2^{2})$ are represented by the green triangle, the released statistics $(\bar{Y}^{*}=-1.3,S^{2}=4.8^{2})$ are represented by the red circle, and the analyst’s posterior mode is represented by the blue diamond. The left and right panels provide the posterior when the statistics are constrained and when the likelihood is constrained, respectively. The shaded area represents the feasible region for $(\mu,\sigma^{2})$ from Theorem 2. This plot is based on 25,000 Gibbs iterations, pruned to every 5th draw.
The tendency of the modified sampler of Ju et al. (2022) to favor values of $(\mu,\sigma^{2})$ that generate samples of $Y_{i}$ relatively far from $[a,b]$ can have advantages in situations where in fact the data values are far from the bounds. However, this tendency can be problematic when data are close to the bounds, as we now illustrate. Consider the same dataset with the same privacy budget, except we suppose the data curator assumes a much wider region for the confidential data: all $Y_{i}\in[0,25]$. We keep the observed statistics as $\bar{Y}=0.5$ and $S^{2}=0.2^{2}$, which are now very close to the boundary of the feasible region.
Figure 11 plots the posterior distributions for the same draw from the Laplace mechanism as in Figure 10 (with updated scale parameter). Our proposed Gibbs sampler from Section 2.2 yields a posterior distribution with its mode near the boundary. Since $\bar{Y}^{*}$ is negative, it focuses the probability density in the part of the feasible region closest to the released value, with substantial uncertainty. However, the modified sampler of Ju et al. (2022) still yields a posterior distribution with substantial probability density near the center of the feasible region. The posterior distribution places fairly little probability density in the region where $\mu$ is near zero, despite this being the most likely region for $\bar{Y}$ under the Laplace Mechanism when the released $\bar{Y}^{*}$ is negative. As discussed above, the need to sample $Y_{i}$ with high density has the strong effect of favoring $(\mu,\sigma^{2})$ values that make it less likely to generate $Y_{i}$ near the boundaries. Additionally, as discussed in Section 4.1, the constrained uniform prior on $(\mu,\sigma^{2})$ induces a marginal prior for $\mu$ with more probability mass in the center of the feasible region, heightening the discrepancy between the observed $(\bar{Y}^{*},S^{2*})$ and the posterior inference for $(\mu,\sigma^{2})$.
I.3 Coverage and Error Comparison
Of course, the previous section considered only two examples to provide intuition. We now compare the behaviors under repeated sampling of our proposed sampler from Section 2.2 and the modified sampler of Ju et al. (2022) described in Section I.1. For this comparison, we use the simulation study described in Section 4.1; the results are presented in Figure 12. Once $n$ is large enough (e.g., $n\geq 100$), we see little difference in the average interval lengths or RMSEs of the point estimators for the two methods. However, until $n$ reaches $500$, the confidence interval coverage rate for the modified sampler of Ju et al. (2022) is notably lower than 95% when the parameters are near the boundary. In contrast, the coverage rates for our sampler are nearly 95% for all values of $n\geq 100$, with similar average lengths and RMSEs as the modified sampler of Ju et al. (2022).
When $n<100$, we see sharper differences in the performances of the two methods. The biggest difference is with respect to the coverage rates. When the true $\mu$ is near the center, both methods have near 100% coverage for small $n$. However, when the true $\mu$ is near the boundary, the likelihood constraint produces interval estimates with very poor coverage for small $n$. The coverage is below 70% for $n<100$ and near zero when $n$ is very small. Our sampler offers higher coverage rates, although they remain below nominal when $n<100$. We also see that, for the smaller values of $n$, the modified sampler of Ju et al. (2022) has shorter average interval lengths than our approach, reflecting its tendency to favor small values of $\sigma^{2}$ as discussed in Section I.2. Additionally, for $\mu$ near the boundary of the parameter space and small $n$, the modified sampler of Ju et al. (2022) tends to have larger RMSE, reflecting the tendency of the sampler (and the prior) to favor central values of $\mu$ as discussed in Section I.2. This natural tendency is a benefit when $\mu=0.5$, but a deficiency when $\mu=0.1$.
In practice, of course, an analyst does not know whether $\mu$ is near the center of the region or near the boundary. One possibility is to spend some privacy budget to assess which model may be preferred. For example, data curators may provide some differentially private measure of how many observations are close to the bounds, which analysts might be able to use to decide which of the models is likely to be more accurate. Additionally, analysts might consider utilizing a different default prior distribution, such as one where the marginal distribution for $\mu$ is uniform. Development of these ideas is an interesting avenue for future work.
Figure 12: Average 95% HPD interval length for $\mu$ (top), root mean square error for estimating $\mu$ (middle), and empirical coverage rate of 95% HPD intervals for $\mu$ (bottom) for different sample sizes. Results based on 10,000 simulated datasets $Y_{i}\in[0,1]$ released with $\varepsilon_{1}=\varepsilon_{2}=1$ and analyzed with prior $p(\mu,\sigma^{2})\propto 1$. Analyses with the statistic constraints from Section 3 are solid lines and with the likelihood constraints from Section I.1 are dotted lines. Data generating model is either ${\mathcal{N}}(\mu=0.1,\sigma^{2}=0.04^{2})$ (red) or ${\mathcal{N}}(\mu=0.5,\sigma^{2}=0.2^{2})$ (blue). Each Gibbs sampler is run for 20,000 iterations.
Appendix J CONSTRAINTS IN SIMPLE LINEAR REGRESSION
@@ I.3 Coverage and Error Comparison
In this section, we derive constraints on the model proposed by Bernstein & Sheldon (2019) in the simple linear regression setting. Letting ${\mathbf{X}}=[{\mathbf{1}}_{n}\hskip 5.78172pt{\mathbf{x}}_{1}]$, the authors release ${\mathbf{X}}^{\top}{\mathbf{X}}$, ${\mathbf{X}}^{\top}{\mathbf{Y}}$, and ${\mathbf{Y}}^{\top}{\mathbf{Y}}$, which include the following elements.
${\mathbf{X}}^{\top}{\mathbf{X}}=\begin{bmatrix}{\mathbf{1}}_{n}^{\top}{\mathbf{1}}_{n}&{\mathbf{1}}_{n}^{\top}{\mathbf{x}}_{1}\\ {\mathbf{x}}_{1}^{\top}{\mathbf{1}}_{n}&{\mathbf{x}}_{1}^{\top}{\mathbf{x}}_{1}\end{bmatrix},\qquad{\mathbf{X}}^{\top}{\mathbf{Y}}=\begin{bmatrix}{\mathbf{1}}_{n}^{\top}{\mathbf{Y}}\\ {\mathbf{x}}_{1}^{\top}{\mathbf{Y}}\end{bmatrix}.$ (107)
@@
    Since $x_{i}\leq 1$ for all $i$, it follows that ${\mathbf{x}}_{1}^{\top}{\mathbf{1}}_{n}={\mathbf{1}}_{n}^{\top}{\mathbf{x}}_{1}=\sum_{i=1}^{n}x_{i}\leq n$.
@@
Appendix K COMPUTE DETAILS
In this section, we describe information on the computer resources needed to reproduce the experiments. We note that Figures 1, 3, 4, 6, 10, and 11 can be run on a personal computer in a few seconds and do not require any additional computing resources. Figure 5 can be run on a personal computer in under five minutes and so also does not require any additional computing resources.